A light truck with manual transmission has a circular brake drum with a diameter of 319 millimeters. Each brake pad, which presses against the drum, is 308 millimeters long. What central angel is subtended by one of the brake pads? Write your answer in both radians and degrees.
Answer by Alan3354(69443)

A light truck with manual transmission has a circular brake drum with a diameter of 319 millimeters. Each brake pad, which presses against the drum, is 308 millimeters long. What central angel is subtended by one of the brake pads? Write your answer in both radians and degrees.
====================
Angle = arc/r = 308/159.5 = 1.931 radians
==============
You can convert to degs